OpenStudy (rowanmartinfan):

Rachna has nickels, dimes, and quarters in her piggy bank. She has $4.65 altogether. She has three more quarters than dimes and twice as many nickels as quarters. How many quarters does Rachna have? A- 8 , B- 11 , C- 22 , D- 41

OpenStudy (anonymous):

Lets the number of quarters is n. What is the number of dimes in terms of n ?

OpenStudy (anonymous):

That is the value of nickels. I am asking the number of dimes in terms of n?

OpenStudy (anonymous):

dimes are 10 cents.

OpenStudy (anonymous):

OK..you have 3 more quarters (n) than dimes that is dimes are n-3 and number of nickels = n/2 So, n*25+(n-3)*10+n/2 * 5 = 465 cents 25n+10n-30+5n/2 = 465 Solve for n.

OpenStudy (anonymous):

You have done wrong calculation. Show me your steps.

OpenStudy (anonymous):

No. But you are close.

OpenStudy (anonymous):

Redo the calculation.

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Show me your steps so that I can correct you..

OpenStudy (anonymous):

You cannot combine 35x and 30. They are different things.

OpenStudy (anonymous):

You have 35x+5x/2 = 465 +30

OpenStudy (anonymous):

35x+2.5x = 495 37.5x = 495

OpenStudy (anonymous):

Solve for x.

OpenStudy (anonymous):

Yes. But that should not be the answer as number of quarters cannot be a fraction!

OpenStudy (anonymous):

Oops.."twice as many nickels as quarters" means nickels are 2n. So n*25+(n-3)*10+2n* 5 = 465 cents Now redo the calculation.

OpenStudy (anonymous):

@RowanMartinFan What is the answer now ?

OpenStudy (anonymous):

25n+10n+10n = 465+30

OpenStudy (anonymous):

45n = 495

OpenStudy (anonymous):

Correct!

OpenStudy (rowanmartinfan):

ok thnak u sir i see now math is bull & i hate it

OpenStudy (anonymous):

Math works on the principal of practice, practice and practice! :)

OpenStudy (rowanmartinfan):

well with my problem my mom tells me i have ADHD & parts of FAS so I cant remember things or stay on task enough to learn things like math @ all